OpenStudy (anonymous):

the sum of the page number on the faceing pages of a book is 89. what are the page numbers? left page= right page=

OpenStudy (anonymous):

Left= 44 Right= 45. I just took a guess at this.

OpenStudy (yttrium):

Provide an equation: let x = page at the left x+1 = page at the right Equation then be: x+(x+1) = 89 2x = 88 x = 44 x+1 = 45 Hence, the page at left is 44 and at the right is 45

OpenStudy (anonymous):

Glad to help!

OpenStudy (anonymous):

Yes, even-numbered pages are on the left-hand side.